OpenStudy (stormswan):

WILL FAN AND MEDAL Can someone PLEASE help me with this! I'm really behind in class and I need to catch up before Winter break is over! :(((

OpenStudy (stormswan):

Given the equation 5x − 4 = –2(3x + 2), solve for the variable. Explain each step and justify your process. Megan solved a similar equation below. Is Megan's solution correct? Explain why or why not. 3(2x − 4) = 5x − 1 6x − 12 = 5x − 1 11x − 12 = –1 11x = 11 x = 1

OpenStudy (stormswan):

this is really urgent and really important! I really want to learn how to do these.

OpenStudy (steve816):

eheee, big nono and a X incorrect, she added 5x instead of subtracting it.

OpenStudy (stormswan):

@steve816 could you please explain to me how to do these types of problems?? Then I won't need to bother anyone on Openstudy.

OpenStudy (stormswan):

For example... this one. 5x − 4 = –2(3x + 2)

OpenStudy (steve816):

Use simple arithmetical logic. That is all I'm going to say.

OpenStudy (stormswan):

thing is, I know how to solve those equations that have the x on one side.. but I don't know how to find ONE x from TWO?? :(

OpenStudy (stormswan):

@steve816 please I really need your help! :(

OpenStudy (stormswan):

@Daniellelovee please can ANYONE help. I'm legit about to cry, this isn't funny anymore. I'm for real trying to learn something but no one will help me.

OpenStudy (anonymous):

The worst thing you can do is freak out. You need to take it step by step. If I have 4 = x + 5, how do we solve for x? We must isolate it, right? Get it by itself? And to do that, that would mean we would have to move the 5 over to the other side. But we can't add it, because that would mean that we end up with a 10 on the right. So therefore we subtract the 5. On the right side, the 5 goes away, and the left side becomes -1 (because 4 - 5 is -1). SIMILARLY, we do the same thing for variables. If I have 4x = 2x + 2, I can start by isolating x. To do that, I must do what I did before in a similar fashion. By subtracting both sides by 2x, the left side becomes 2x and the right side just becomes 2. Therefore, 2x = 2. Solving for x by dividing both sides by 2, we get x = 1. This algebraic simplification is paramount to solving the problem that you originally posted.

OpenStudy (dayakar):

5x-4= -2(3x+2) 5x-4= -6×-4 (using distributive law)

OpenStudy (dayakar):

Add both sides 6x simplify it

OpenStudy (dayakar):

R u doing

OpenStudy (nincompoop):

5x − 4 = –2(3x + 2) try to expand the right-hand-side first

OpenStudy (dayakar):

11x-4 =-4

OpenStudy (dayakar):

11x=0

OpenStudy (dayakar):

Then X=0

OpenStudy (stormswan):

@dayakar wow, thank you so much!!! this actually helped. I totally get it now! I didm know I was supposed to use the distributive law to do this. I guess thats where I was stuck at. thank you so much! at least someone helped. 5x-4= -2(3x+2) 5x-4= -6×-4

OpenStudy (stormswan):

@nincompoop @cshrix thanks for trying to help me. I know I shouldn't freak out, but my parents are being too hard on me and I know that if I don't learn this stuff soon, I'll be dead within a few days.(not literally, but you know what I mean)